# The Smith form of Sylvester and Bézout matrices for zero-dimensional ideals

### v1

@@ Abstract
Let $\mathbb{K}$ be a field and let $f,g\in\mathbb{K}[x,y]$ be such that the ideal $\langle f,g\rangle$ is zero-dimensional. We study the Sylvester and Bézout resultant polynomial matrices, built by interpreting $f$ and $g$ as univariate polynomials in $x$ with coefficients in $\mathbb{K}[y]$. We characterize their Smith forms over $\mathbb{K}[y]$ in terms of the dual spaces of differential operators, that were defined and studied by H. M. Möller et al. In particular, we show that, if the leading coefficients of $f$ and $g$ are coprime over $\mathbb{K}[y]$, then the partial multiplicities of the Sylvester and Bézout resultant matrices coincide with certain integers, that we call Möller indices. These indices are uniquely determined by $\langle f,g\rangle$, and can be easily computed from a Gauss basis, as defined in [M. G. Marinari, H. M. Möller, T. Mora, Trans. Amer. Math. Soc. 348(8):3283–3321, 1996], of the dual spaces. We then generalize this result to the case of common factors in the leading coefficients, which correspond to intersections at $x=\infty$, again describing all the invariant factors of Sylvester and Bézout resultant matrices. As a corollary, this fully characterizes the algebraic multiplicity of all the roots of the resultant $\mathrm{Res}_{x}(f,g)\in\mathbb{K}[y]$ in terms of the intersection multiplicities for $f$ and $g$, including those arising from infinite intersections. We discuss both algebraic and computational implications of our results.
@@ 2.2 Basic properties of Sylvester and Bézout matrices
$\Lambda_{m}(z)^{\top}B_{f,g}\Lambda_{n}(x)=\mathcal{B}_{f,g}(x,z)=\frac{f(x)g(z)-f(z)g(x)}{x-z}.$
@@
Let $S$ be the $(k,k)$-Sylvester matrix of $f$ and $g$. Proposition 2.10 and Sylvester’s rank inequality yield $2\mathrm{rank}(S)-2k\leq 2~\mathrm{rank}(B),$ implying in turn $\dim\ker B\leq\dim\ker S.$ In view of this observation, it suffices to verify $B\Lambda_{k}^{(i)}(x_{0})=0$ for each allowed choice of $x_{0}$ and $i$, since (applying Lemma 2.13 in the case $m=n=k$ and hence $s=0$) these equations also imply $\dim\ker B\geq\dim\ker S$.
@@ 3 Möller indices
    $G_{i}=\mathrm{LM}(G_{i})+\psi_{i},\ \psi_{i}\in V$, and for all $i,j$, $\mathrm{LM}(G_{j})$ is not a monomial in $\psi_{i}$.
@@
Figure 1: Visualization of Example 3.12 by Young diagrams with width $\beta$. Left: Leading monomials, French style diagram. The $i$-th column corresponds to the subset $\mathcal{L}_{i-1}$ in the partition of $\mathrm{LM}(V)$, and thus its height is $\alpha(i)$. Moving leftwards (resp. upwards) correspond to the action of $A_{x}$ (resp. $A_{y}$). Right: Corresponding vectors in the Gauss basis $\{G_{i}\}_{i=1}^{6}$, English style diagram. The height of each column still captures the Möller indices with respect to $y$. The leading vectors ($\phi_{0}=G_{6}$, $\phi_{1}=G_{4}$) appear on top of each column.
@@ 4.1 Case of coprime leading coefficients
Let $a\in R[x]_{m+n-1}$, be such that $\phi(a)=[0]_{N}$, that is, $a\in I$. Since the leading coefficient of $f$ is a unit of $R$, by Lemma 4.1 there exist $u\in R[x]_{n}$ and $v\in R[x]_{m}$ such that $a=uf+vg$. Hence, $a\in\mathrm{Im}\ \mathcal{S}$ and $a]_{M}=[0]_{M}$.
@@ 4.3 Smith form of $S(y)$ when there are roots at infinity
    There is a multiplicity-preserving bijection between the finite roots $(x_{0},y_{0})$ of $f$ and $g$ and $(\phi^{-1}(x_{0}),y_{0})$ of $\hat{f}$ and $\hat{g}$. Furthermore, the multiplicity of the root $(\infty,y_{0})$ for $f$ and $g$ is equal to the multiplicity of the root $(d/c,y_{0})$ of $\hat{f}$ and $\hat{g}$.
@@
    If $(\infty,y_{0})$ is an infinite intersection of $f$ and $g$, let $x_{i}\in\mathbb{F}$, $i=1,\dots,\ell$, denote all the elements of $\mathbb{F}$ such that $f(x_{i},y_{0})=g(x_{i},y_{0})=0$, and let $\alpha_{i}(0),\dots,\alpha_{i}(\beta_{i}-1)$ be the Möller indices with respect to $y$ of the the ideal $\langle f,g\rangle$ at the point $P_{i}=(x_{i},y_{0})$. Moreover, let $\alpha_{\infty}(0),\dots,\alpha_{\infty}(\beta_{\infty}-1)$ be the Möller indices with respect to $y$ of $f$ and $g$ at $(\infty,0)$. Then, the partial multiplicities of $y_{0}$ in $S(y)$ are
@@ 5 Smith form of Bézout matrices
Let $B(y)\in\mathbb{F}[y]^{k\times k}$ be the Bezóut matrix of $f,g\in\mathbb{F}[x,y]$, where $k=\max\{\deg_{x}f,\deg_{x}g\}$ and suppose $\langle f,g\rangle$ is zero-dimensional. Fix $y_{0}\in\mathbb{F}$. Let $P_{i}=(x_{i},y_{0})$, $i=1,\dots,\ell$, where either $x_{i}\in\mathbb{F}$ if $(x_{i},y_{0})\in\mathcal{V}(I)$ or $x_{i}=\infty$ if $(\infty,y_{0})$ is an infinite intersection of $f$ and $g$, according to the modified version of Definition 4.13 as described in Remark 4.19. Then, the partial multiplicities of $y_{0}$ in $B(y)$ are

### v2

@@ Abstract
Let $\mathbb{K}$ be a field and let $f,g\in\mathbb{K}[x,y]$ be such that the ideal $\langle f,g\rangle$ is zero-dimensional. We study the Sylvester and Bézout resultant polynomial matrices, built by interpreting $f$ and $g$ as univariate polynomials in $x$ with coefficients in $\mathbb{K}[y]$. We characterize their Smith forms over $\mathbb{K}[y]$ in terms of the dual spaces of differential operators, that were defined and studied by H. M. Möller et al. In particular, if $\mathbb{K}$ is algebraically closed we show that, if the leading coefficients of $f$ and $g$ are coprime over $\mathbb{K}[y]$, then the partial multiplicities of the Sylvester and Bézout resultant matrices coincide with certain integers, that we call Möller indices. These indices are uniquely determined by $\langle f,g\rangle$, and can be easily computed from a Gauss basis, as defined in [M. G. Marinari, H. M. Möller, T. Mora, Trans. Amer. Math. Soc. 348(8):3283–3321, 1996], of the dual spaces. We then generalize this result to the case of common factors in the leading coefficients, which correspond to intersections at $x=\infty$, again describing all the invariant factors of Sylvester and Bézout resultant matrices. As a corollary, this fully characterizes the algebraic multiplicity of all the roots of the resultant $\mathrm{Res}_{x}(f,g)\in\mathbb{K}[y]$ in terms of the intersection multiplicities for $f$ and $g$, including those arising from infinite intersections. We discuss both algebraic and computational implications of our results.
@@ 2.2 Basic properties of Sylvester and Bézout matrices
$\Lambda_{k}(z)^{\top}B_{f,g}\Lambda_{k}(x)=\mathcal{B}_{f,g}(x,z)=\frac{f(x)g(z)-f(z)g(x)}{x-z}.$
@@
Let $S$ be the $(k,k)$-Sylvester matrix of $f$ and $g$. Proposition 2.10 and Frobenius’s rank inequality yield $2\mathrm{rank}(S)-2k\leq 2~\mathrm{rank}(B),$ implying in turn $\dim\ker B\leq\dim\ker S.$ In view of this observation, it suffices to verify $B\Lambda_{k}^{(i)}(x_{0})=0$ for each allowed choice of $x_{0}$ and $i$, since (applying Lemma 2.13 in the case $m=n=k$ and hence $s=0$) these equations also imply $\dim\ker B\geq\dim\ker S$.
@@ 3 Möller indices
    $G_{i}=\mathrm{LM}(G_{i})+\psi_{i},\ \psi_{i}\in\mathcal{D}$, and for all $i,j$, $\mathrm{LM}(G_{j})$ is not a monomial in $\psi_{i}$.
@@
Figure 1: Visualization of Example 3.12 by Young diagrams with width $\beta$. Left: Leading monomials, French style diagram. The $i$-th column corresponds to the subset $\mathcal{L}_{i-1}$ in the partition of $\mathrm{LM}(V)$, and thus its height is $\alpha(i)$. Moving leftwards (resp. downwards) correspond to the action of $A_{x}$ (resp. $A_{y}$). Right: Corresponding vectors in the Gauss basis $\{G_{i}\}_{i=1}^{6}$, English style diagram. The height of each column still captures the Möller indices with respect to $y$. The leading vectors ($\phi_{0}=G_{6}$, $\phi_{1}=G_{4}$) appear on top of each column.
@@ 4.1 Case of coprime leading coefficients
Let $a\in R[x]_{m+n-1}$, be such that $\phi(a)=[0]_{N}$, that is, $a\in I$. Since the leading coefficient of $f$ is a unit of $R$, by Lemma 4.1 there exist $u\in R[x]_{n-1}$ and $v\in R[x]_{m-1}$ such that $a=uf+vg$. Hence, $a\in\mathrm{Im}\ \mathcal{S}$ and $a]_{M}=[0]_{M}$.
@@ 4.3 Smith form of $S(y)$ when there are roots at infinity
    There is a multiplicity-preserving bijection between the finite roots $(x_{0},y_{0})$ of $f$ and $g$ and $(\phi^{-1}(x_{0}),y_{0})$ of $\hat{f}$ and $\hat{g}$. Furthermore, the multiplicity of the root $(\infty,y_{0})$ for $f$ and $g$ is equal to the multiplicity of the root $(-d/c,y_{0})$ of $\hat{f}$ and $\hat{g}$.
@@
    If $(\infty,y_{0})$ is an infinite intersection of $f$ and $g$, let $x_{i}\in\mathbb{F}$, $i=1,\dots,\ell$, denote all the elements of $\mathbb{F}$ such that $f(x_{i},y_{0})=g(x_{i},y_{0})=0$, and let $\alpha_{i}(0),\dots,\alpha_{i}(\beta_{i}-1)$ be the Möller indices with respect to $y$ of the the ideal $\langle f,g\rangle$ at the point $P_{i}=(x_{i},y_{0})$. Moreover, let $\alpha_{\infty}(0),\dots,\alpha_{\infty}(\beta_{\infty}-1)$ be the Möller indices with respect to $y$ of $f$ and $g$ at $(\infty,y_{0})$. Then, the partial multiplicities of $y_{0}$ in $S(y)$ are
@@ 5 Smith form of Bézout matrices
Let $B(y)\in\mathbb{F}[y]^{k\times k}$ be the Bezóut matrix of $f,g\in\mathbb{F}[x,y]$, where $k=\max\{\deg_{x}f,\deg_{x}g\}$ and suppose $\langle f,g\rangle$ is zero-dimensional. Fix $y_{0}\in\mathbb{F}$. Let $P_{i}=(x_{i},y_{0})$, $i=1,\dots,\ell$, where either $x_{i}\in\mathbb{F}$ if $(x_{i},y_{0})\in\mathcal{V}(I)$ or $x_{i}=\infty$ if $(\infty,y_{0})$ is an infinite intersection of $f$ and $g$ defined with respect to the coefficients of $x^{k}$ in both $f$ and $g$, i.e., according to the modified version of Definition 4.13 as described in Remark 4.19. Then, the partial multiplicities of $y_{0}$ in $B(y)$ are